OpenStudy (anonymous):

solve 1/4(5x- 1/2) -5/3<1/3

OpenStudy (anonymous):

write it with equation forum

OpenStudy (anonymous):

do you mean\[\frac{1}{4}(5x-\frac{1}{2})-\frac{5}{3}<\frac{1}{3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do I times the out side by the numbers inside ()?

OpenStudy (anonymous):

i hate writing through eq. editor

OpenStudy (anonymous):

mutliply through by LCD 12 to clear denominators \[3(5x-1/2)-20<4\]\[15x-3/2-20<4\] you can multply through by 2 if you want to clear the last denominator \[30x-3-40<8\]\[30x<51\]\[x<51/30\]I hope I didn't make an arithmetic error!!!!!

OpenStudy (anonymous):

instead of distribution, add 20 to both sides\[3(5x-1/2)<24\]then divide through by 3\[5x-1/2<8\]now add 1/2 to both sides\[5x<\frac{17}{2}\]now multiply through by the reciprocal of 5\[x<\frac{17}{2}*\frac{1}{5}\]\[x<\frac{17}{10}\]which is the rduced form of 51/30 Maybe i didn't make any errors!!!